# Foundation Cures Personalization: Improving Personalized Models’ Prompt Consistency via Hidden Foundation Knowledge

### v2

@@ Abstract
Facial personalization faces challenges to maintain identity fidelity without disrupting the foundation model’s prompt consistency. The mainstream personalization models employ identity embedding to integrate identity information within the cross-attention mechanisms of UNet. However, our preliminary experimental findings reveal that identity embeddings compromise the effectiveness of other tokens in the prompt, thereby limiting high prompt consistency and controllability. Moreover, by deactivating identity embedding, personalization models still demonstrate the underlying foundation models’ ability to control facial attributes precisely. It suggests that such foundation models’ knowledge can be leveraged to cure the ill-aligned prompt consistency of personalization models. Building upon these insights, we propose FreeCure, a framework that improves the prompt consistency of personalization models with their latent foundation models’ knowledge. First, by setting a dual inference paradigm with/without identity embedding, we identify attributes (e.g., hair, accessories, etc.) for enhancements. Second, we introduce a novel foundation-aware self-attention module, coupled with an inversion-based process to bring well-aligned attribute information to the personalization process. Our approach is training-free, and can effectively enhance a wide array of facial attributes in a non-intrusive manner; and it can be seamlessly integrated into existing popular personalization models, without harming their well-trained modules. FreeCure has demonstrated significant improvements in prompt consistency across a diverse set of state-of-the-art facial personalization models while maintaining the integrity of original identity fidelity. The project page is available here.
@@
1 Introduction
Human face-centric personalization represents compelling downstream applications of art creation, advertising, and entertainment in the realm of text-to-image synthesis  [17, 40]. Given a limited number of images that depict particular identities, facial personalization models generate novel content that reflects these identities through user-defined text prompts  [11, 35]. However, this aspiration is hindered by a persistent challenge: the necessity to maintain high fidelity to the identity while ensuring the controllability of the generated content, also referred to as prompt consistency  [53, 20]. This challenge is pronounced in facial personalization since any imperfections or misalignment in the generated faces are particularly salient to humans. Therefore, human face-centric personalization mandates dedicated attention and research efforts.
Previous research of facial personalization aims to integrate identity-specific information into the cross-attention mechanisms, with either fine-tuning based strategy  [35, 11] or tuning-free paradigm with an identity encoder  [46, 47, 27, 14]. However, the dual objectives of maintaining prompt consistency and identity fidelity remain unresolved. Despite the aforementioned advancements, state-of-the-art personalization techniques still struggle to enhance identity fidelity without sacrificing prompt consistency.
Our preliminary experiments on a wide range of personalization methods indicate that, under identical experimental conditions (e.g., prompts and initial random noise), personalization models’ outputs exhibit a significant decline in prompt consistency compared to generated results without identity embedding. For instance, as shown in the first two examples in Figure 1, personalization models fail to generate “blonde curly hair” and “angry” faces accurately, whereas their counterparts without identity embedding can handle these prompts in a highly faithful manner. Plus, when handling more complex prompts that consist of multiple attributes, the personalization model generates even poorer results. These findings show a fact overlooked by most previous works: while it is true that personalization models show their degradation in prompt consistency, the ability of their original foundation models is unharmed but overridden. Based on this observation, we step further to focus on the most critical components in personalization models: identity embedding and attention mechanisms. We find that identity embeddings can significantly undermine prompt consistency by impeding normal representation of other attribute-related tokens within the prompt through cross-attention mechanisms. This, in turn, adversely affects their effective expression in the latent space of the U-Net. However, given the compelling zero-shot identity extraction capability of identity embedding, directly modifying personalization models’ well-trained cross-attention modules can destroy their ability to capture precise identity information. Therefore, our core pursuit is clear: Is it feasible to mitigate the erosion of prompt consistency in personalization models while keeping their trained cross-attention modules unaffected?
@@ 1 Introduction
  • We propose a training-free framework that leverages high prompt consistency information from foundation models to enhance multiple weakened attributes in the synthesized results of personalization models. Crucially, the enhancement of various attributes is achieved without mutual interference.
  • Our framework can be seamlessly integrated into widely used personalization models and diverse foundational models. Comprehensive experiments show that our approach improves prompt consistency while maintaining the well-trained ability for identity preservation.
@@ 2 Related Work
Identity-preserving Generation. Identity-preserving generation can be broadly categorized into fine-tuning-based and encoder-based approaches. Fine-tuning-based methods  [11, 43, 1, 54, 10, 50, 31] either optimize a vector within a textual embedding to encode identity, modify specific weights in the U-Net [35, 23, 5, 19, 13, 36] or use low-rank decomposition techniques [18]. However, these methods require training a separate model for each identity, which limits their scalability and poses a risk of overfitting, particularly when working with limited data. In contrast, encoder-based methods [46, 48, 29, 27, 32, 7, 38, 26] focus on training an image encoder or a lightweight mapping network to derive identity representations aligning with textual embeddings. This allows for zero-shot personalization with novel identity references. Additionally, these methods also integrate key and value components of cross-attention maps into the training process [47, 32, 29, 44, 42, 39]. By employing detection and segmentation models [8, 21], they refine identity alignment and generation diversity by eliminating extraneous background elements. [14] integrates a lightning-fast T2I model to enable rapid sampling during personalization. Overall, encoder-based methods offer more flexible interactions between identity and textual embeddings but they still struggle to balance prompt consistency and identity fidelity.
Attention in Stable Diffusion. The attention mechanism is a fundamental element in Stable Diffusion.  [15, 2, 12, 22, 41, 4, 30, 9] have made attempts to edit important attention layers within diffusion models, allowing for localized attribute editing and style transfer.  [28] has explored the distinct roles of self-attention and cross-attention layers within the U-Net.  [24, 6, 3, 55] have investigated the latent properties of diffusion models to identify an editable semantic space via latent space analysis. With the advancement of inversion techniques  [40] in diffusion models, these methods aim to alter specific attributes of realistic images. Overall, the attention mechanism provides valuable insights into diffusion models. Nevertheless, within the realm of personalization, the role of identity embedding within the attention mechanism remains under-explored.
@@ 3 Revisit ID Embedding in Personalization
We conduct a comprehensive analysis to justify the limitations of current personalization methods in maintaining prompt consistency and to uncover the underlying factors contributing to this issue. This investigation not only highlights the challenges inherent in existing approaches but also provides a foundational basis and critical insights that inform the development of our proposed methodology.
Figure 2: Limitation of identity embedding in personalization models. (a) compares the impact of identity embeddings on the personalization and foundation outputs (zoom-in for better evaluation of prompt consistency). (b)-(d) compare scores of the cross-attention maps corresponding to identity embeddings and other attribute-related token embeddings.
@@ 3.1 Dual Denoising to Study Prompt Consistency
To elucidate current personalization methods’ limitations, we have devised a simple comparative experiment as an initial exploration. We maintain a fixed noisy latent code $z_{T}$ and implement two parallel denoising procedures. The only difference is that one denoising procedure incorporates textual embeddings fused with identity embedding $\tilde{c}$, while the other uses the original textual embeddings $c$, without any identity embedding information. For clarity, we will refer to the denoising procedure with identity embedding (personalization denoising) as PD, and the denoising process that excludes identity embedding (foundation denoising) as FD hereafter. These two procedures can be represented as the following equations:
$\textbf{PD}:\epsilon_{p}=\epsilon_{\theta}(z_{t},t,\tilde{c}),$ (1)
$\textbf{FD}:\epsilon_{f}=\epsilon_{\theta}(z_{t},t,c),$ (2)
where $\epsilon_{p}$ and $\epsilon_{f}$ denote the predicted noise from PD and FD, respectively. We have conducted these experiments using two facial personalization methods  [47, 27] and results are presented in Column (a) of Figure 2. It is evident that the output from PD exhibits a reduced ability to match the specific facial attributes compared to the counterpart without identity embedding. For instance: 1) The hair features do not align with the prompt’s specified “black curly” as expected; 2) The expression “laughing” is also restricted. Conversely, the results from FD show greater prompt consistency. These findings indicate that current identity embedding methods may not effectively address “the balance between identity fidelity and prompt consistency”.
@@ 3.2 ID Embedding’s Effect on Attention Layers
OBSERVATION 1: ID embeddings disrupt cross-attention layers. To investigate the underlying causes of this phenomenon, we conduct a visualization experiment utilizing the technique of cross-attention map analysis  [15, 2], because identity embedding is primarily fused into cross-attention layers. Columns (b-d) of Figure 2 present the visualization results of cross-attention maps for both PD and FD processes, focusing on identity embeddings and the attribute-specific token embeddings for “black curly hair” and “laughing”. It reveals that the identity embedding significantly amplifies activation in facial regions while simultaneously reducing activation for other tokens and disrupting their typical interactions within the cross-attention mechanism. This finding highlights identity embedding as a pivotal factor contributing to the reduced prompt consistency observed in personalization models. Additionally, we observe that during the FD process, where identity embedding injection is absent, prompt consistency remains preserved. This observation underscores a critical insight: although most personalization models are well fine-tuned, their inherent capability to generate faces with high prompt consistency is preserved but unexpectedly overridden by external identity embeddings. Thus, it is reasonable to conclude that the FD procedure functions in a manner analogous to pretrained foundation models  [34, 33], as implied by its full name, “Foundation Denoising”, and foundation models’ knowledge is still maintained in personalization models.
OBSERVATION 2: Cross-attention layers are highly susceptible. We conduct another experiment on cross-attention layers to prove its susceptibility, which involves incorporating a portion of the cross-attention map from FD’s placeholder token into the cross-attention map of PD’s identity embedding. The formula for this approach is
@@
where $A^{p}_{:,:,m}$ represents the cross-attention map of PD’s identity embedding, and $A^{f}_{:,:,n}$ represents the cross-attention map of the FD’s placeholder embedding (“man”, “woman”, etc.), whose positions in their respective prompts are $m$ and $n$. The parameter $\alpha$ controls the weight of the FD’s cross-attention map that is injected. Spatial indices $i$ and $j$ of $A$ are omitted since the interpolation operation is performed across the entire attention map. We apply this “interpolation” process to the same examples discussed in Figure 2 and the results are shown in Figure. 3. It is observed that, as the weight of the map from the FD process increases, the model incrementally loses identity fidelity, even though it regains the facial attributes that were missing before. This experiment further demonstrates that the cross-attention mechanisms in personalization models are rather susceptible to modifications. To preserve the models’ capacity for identity preservation, it is better to leave these well-trained cross-attention modules unaltered.
Figure 3: Interpolation of identity embedding’s cross-attention map. By injecting a portion $\alpha$ of the cross-attention map from the FD to the PD process, attributes can be restored. However, the identity fidelity can be easily undermined.
Figure 4: The overview of FreeCure. For a trained personalization model $\epsilon_{\theta}$, we first introduce (a): dual inference paradigm to generate faces with/without identity ($I_{p}$ and $I_{f}$), where $I_{f}$ without identity embedding shows better prompt consistency. Next, we leverage an external segmentation model $\Psi(\cdot)$ to derive related masks of target attributes with clear spatial information (hair, sunglasses, etc.) and merge them into a mask $\mathcal{M}$. In (b): we modify the original self-attention modules in the U-Net with our proposed FASA (c), which concatenates key and value matrices of FD process and PD process, together with a scaling mask to achieve the attribute injection. Finally, we utilize a simple yet effective strategy (d): asymmetric prompt guidance (APG) to restore abstract attributes (e.g., expressions).
OBSERVATION 3: Self-attention layers are the key to “foundation knowledge”. Our preliminary findings demonstrate that identity embeddings in cross-attention layers are disrupting personalization models’ prompt consistency. In contrast, its strong ability of identity extraction makes it rather challenging to be modified. Given this “dilemma”, we aim to explore a new approach to enhance prompt consistency by leveraging the self-attention layers in the U-Net. Since most personalization models introduce minimal modifications to self-attention layers, it is reasonable to assume that the aforementioned hidden foundation knowledge is preserved within them. By enhancing the self-attention layers in personalization models while keeping cross-attention layers intact, we anticipate achieving better alignment of facial attributes in personalized generation, all while minimizing adverse effects on identity fidelity.
@@ 4 Methodology
Building upon our investigation in Section 3, we propose FreeCure, a training-free framework designed to improve the prompt consistency of facial personalization models, particularly in contexts involving multiple attributes. We classify facial attributes into two distinct categories: those with well-defined spatial localization (e.g., hair, accessories) and those that are more abstract in nature (e.g., expression), implementing different strategies to enhance each category. Initially, we develop a foundation-aware self-attention (FASA) mechanism to integrate localized attributes from the FD process into the PD process (Section 4.1). Subsequently, we apply an asymmetric prompt guidance (APG) strategy to reconstruct more abstract attributes (Section 4.2). The overview is illustrated in Figure 4.
@@ 4.1 Foundation-Aware Self-Attention
Obtain Masks of Spatial Localized Attributes. As shown in Figure 4 (a), we adapt the dual inference branches which are identical to the FD/PD process introduced in Section 3 for personalization models to generate the personalized face $I_{p}$ with unsatisfactory prompt consistency as well as the foundation face $I_{f}$ with high prompt consistency. Next, we utilize an external face parsing model, denoted as $\Psi(\cdot)$, to extract the binary mask $M_{i}$ corresponding to $A_{i}$ from the foundation outputs $I_{f}$. Generally, when restoring multiple attributes, we simply need to compute the different masks respectively and merge them =$\mathcal{M}=\bigcup\{M_{i}\}$. $\mathcal{M}$ contains the spatial information of attributes that align with the target prompt, which will play an important role in the attribute restoration process that will be mentioned in the next part.
FASA Mechanism. FASA is the core module that links the information between FD and PD processes and enables personalization models to restore the attributes via their foundation knowledge. As shown in Figure 4(c), we at first identify the self-attention modules in the model. Specifically, in each timestep $t$ and attention layer $l$, we denote PD and FD’s key, query, and value matrices as $\mathcal{KQV}^{tl}_{p}=\{K^{tl}_{p},Q^{tl}_{p},V^{tl}_{p}\}$ and $\mathcal{KQV}^{tl}_{f}=\{K^{tl}_{f},Q^{tl}_{f},V^{tl}_{f}\}$. Second, we concatenate the key and value matrices of FD process to those of PD process: $\hat{K^{tl}}=[K^{tl}_{p},K^{tl}_{f}]$, $\hat{V^{tl}}=[V^{tl}_{p},V^{tl}_{f}]$. Thus, by omitting the labels of $t$ and $l$ for simplicity, the operation of FASA is
@@
Fine-grained Restoration with Scaling Masks. The approach described in Equation 4 allows the PD process to obtain information from the FD process. However, we observe that the method also retains a substantial degree of unrelated features, leading to a notable loss of identity fidelity, as shown in Figure 5. To constrain attribute restoration in a specific region without disrupting identity information, we apply pre-calculated masks $\mathcal{M}$ of different attributes to the similarity map. Additionally, to further enhance FASA’s performance, we introduce an additional scaling factor, denoted as $\omega$, to control the magnitude of injecting attribute features from FD to PD. Therefore, the enhanced FASA mechanism can be written as
@@
Where $\mathbf{1}$ denotes a matrix with all elements equal to 1, designed to preserve attention on the original features within the PD process. $\odot$ represents the Hadamard product. With the enhanced FASA, attributes of the FD process can be successfully extracted and precisely injected into the PD process without disrupting the original personalization model’s ability of maintaining identity fidelity.

### v3

@@ Abstract
Facial personalization faces challenges to maintain identity fidelity without disrupting the foundation model’s prompt consistency. The mainstream personalization models employ identity embedding to integrate identity information within the attention mechanisms. However, our preliminary findings reveal that identity embeddings compromise the effectiveness of other tokens in the prompt, thereby limiting high prompt consistency and attribute-level controllability. Moreover, by deactivating identity embedding, personalization models still demonstrate the underlying foundation models’ ability to control facial attributes precisely. It suggests that such foundation models’ knowledge can be leveraged to cure the ill-aligned prompt consistency of personalization models. Building upon these insights, we propose FreeCure, a framework that improves the prompt consistency of personalization models with their latent foundation models’ knowledge. First, by setting a dual inference paradigm with/without identity embedding, we identify attributes (e.g., hair, accessories, etc.) for enhancements. Second, we introduce a novel foundation-aware self-attention module, coupled with an inversion-based process to bring well-aligned attribute information to the personalization process. Our approach is training-free, and can effectively enhance a wide array of facial attributes; and it can be seamlessly integrated into existing popular personalization models based on both Stable Diffusion and FLUX. FreeCure has consistently shown significant improvements in prompt consistency across these facial personalization models while maintaining the integrity of their original identity fidelity.
1 Introduction
Human face-centric personalization represents compelling downstream applications of art creation, advertising, and entertainment in the realm of text-to-image synthesis  ho2020denoising ; song2020denoising . Given a limited number of images that depict particular identities, facial personalization models generate novel content that reflects these identities through diverse conditions  gal2023an ; ruiz2023dreambooth ; zhang2023adding ; mou2024t2i . However, this aspiration is hindered by a persistent challenge: the necessity to maintain high fidelity to the identity while ensuring the controllability of the generated content, also referred to as prompt consistency  zhang2024survey ; huang2024diffusion . This challenge is pronounced in facial personalization since any imperfections or misalignment in the generated faces are particularly salient to humans. Therefore, compared to common object personalization, human face-centric personalization mandates dedicated attention and research efforts.
Previous research of facial personalization aims to integrate identity-specific information into the cross-attention mechanisms, with either fine-tuning based strategy  ruiz2023dreambooth ; gal2023an or tuning-free paradigm with an identity encoder  xiao2024fastcomposer ; li2024photomaker ; NEURIPS2024_409fcc9d ; wang2024instantid ; jiang2025infiniteyou . However, the dual objectives of maintaining prompt consistency and identity fidelity remain unresolved. Despite the aforementioned advancements, state-of-the-art personalization techniques still struggle to enhance identity fidelity without sacrificing prompt consistency.
@@ 1 Introduction
Our preliminary experiments indicate that, under identical experimental conditions (e.g., prompts and initial random noise), personalization models’ outputs exhibit a significant decline in prompt consistency compared to generated results without identity embedding. For instance, as shown in the first two examples in Fig.1, personalization models fail to generate “blonde curly hair" and “angry" faces accurately, whereas their counterparts without identity embedding can handle these prompts in a highly faithful manner. Plus, when handling more complex prompts that consist of multiple attributes, the personalization model generates even poorer results. These findings show a fact overlooked by most previous works: while personalization models show their degradation in prompt consistency, the ability of their original foundation models is unharmed but overridden. Based on this observation, we further find that identity embeddings can significantly undermine prompt consistency by impeding normal representation of other attribute-related tokens within the prompt through cross-attention mechanisms. This, in turn, adversely affects their effective expression in the latent space of the U-Net. However, given the compelling zero-shot identity extraction capability of identity embedding, directly modifying personalization models’ well-trained cross-attention modules can destroy their ability to capture precise identity information. Therefore, our core pursuit is clear: Is it feasible to mitigate the erosion of prompt consistency in personalization models while keeping their trained cross-attention modules unaffected?
@@
  • We propose a training-free framework that leverages high prompt consistency information from foundation models to enhance multiple weakened attributes generated by personalization models. The enhancement of various attributes is achieved without mutual interference.
  • Our framework can be seamlessly integrated into widely used personalization models and diverse foundational models, including Stable Diffusion and FLUX. Comprehensive experiments show that our approach improves prompt consistency while maintaining the well-trained ability for identity preservation.
@@ 2 Related Work
Identity-preserving Generation. Identity-preserving generation can be broadly categorized into fine-tuning-based and encoder-based approaches. Fine-tuning-based methods  gal2023an ; voynov2023p+ ; alaluf2023neural ; han2023svdiff ; zhang2023prospect ; dong2022dreamartist ; yuan2024inserting ; pang2024cross either optimize a vector within a textual embedding to encode identity, modify specific weights in the model ruiz2023dreambooth ; Kumari_2023_CVPR ; chae2023instructbooth ; hua2023dreamtuner ; arar2023domain ; gu2024mix ; ruiz2024hyperdreambooth ; tewel2023key or use LoRA techniques hu2021lora . However, these methods require training a distinct model for each identity, which compromises scalability and increases susceptibility to overfitting. In contrast, encoder-based methods utilize large-scale pretraining to automatically derive identity representations aligning with textual embeddings, enabling zero-shot personalization with novel identity references. Some methods  wei2023elite ; chen2024dreamidentity ; xiao2024fastcomposer ; shiohara2024face2diffusion ; wang2024high ; li2024photomaker ; peng2024portraitbooth only train a mapping network or modify cross-attention weights to embed identity information, while others  ye2023ip ; gal2024lcm ; wang2025dp ; ma2024subject ; wang2024instantid ; valevski2023face0 ; li2024stylegan ; zhang2024flashface ; cui2024idadapter incorporate cross-attention adapters during training. Recent works  NEURIPS2024_409fcc9d ; jiang2025infiniteyou also leverages Diffusion Transformers (DiT) bflflux ; peebles2023scalable ; 10.5555/3692070.3692573 to reach more impressive performance. In this work, we focus primarily on encoder-based methods, as they represent the state-of-the-art paradigm for personalization and demonstrate particular efficacy in facial performance.
Attention in Diffusion Models. Attention mechanisms serve as foundational components in text-to-image diffusion models. Prior research  hertz2023prompttoprompt ; gu2024photoswap ; Zhao_2023_ICCV ; kong2024omg ; tumanyan2023plug ; cao2023masactrl ; liu2024towards ; kong2024anymaker ; ma2024character ; patashnik2025nested has leveraged semantic information from cross-attention maps to facilitate object-level editing, while other methods  tewel2024training ; alaluf2024cross ; nam2024dreammatcher ; chefer2023attend ; wang2024magicface ; hertz2024style ; ding2024freecustom ; tewel2024add ; tan2024ominicontrol ; wang2024moa have employed spatial features from self-attention layers to achieve more precise modifications or style transfer. However, the positional information derived from attention maps are inherently constrained to object-level manipulation, making them less robust for fine-grained facial attribute generation. Plus, the function of face-centric embeddings within attention mechanisms remains insufficiently explored, especially in facial personalization models.
@@ 3 Revisit ID Embedding in Personalization
We conduct a comprehensive analysis to justify the limitations of current personalization methods in maintaining prompt consistency. This investigation highlights the challenges inherent in existing approaches provides a foundational basis and critical insights for our proposed methodology.
@@ 3.1 Dual Denoising to Study Prompt Consistency
To elucidate current personalization methods’ limitations, we devise a comparative experiment as an initial exploration. We maintain a fixed noisy latent code $z_{T}$ and implement two parallel denoising procedures  brooks2023instructpix2pix . The only difference is that one denoising procedure incorporates textual embeddings $c$ and identity embedding $c_{id}$, while the other set $c_{id}$ into a zero tensor $\tilde{c_{id}}$. For clarity, we will refer to the denoising procedure with identity embedding (personalization denoising) as PD, and the denoising process that excludes identity embedding (foundation denoising) as FD hereafter:
[…]
where $\epsilon_{p}$ and $\epsilon_{f}$ denote the predicted noise from PD and FD, respectively. We have conducted these experiments using two facial personalization methods  xiao2024fastcomposer ; li2024photomaker and results are presented in Column (a) of Fig.2. It is evident that the output from PD exhibits a reduced ability to match the specific facial attributes compared to the counterpart without identity embedding. For instance: 1) The hair features do not align with the prompt’s specified “black curly" as expected; 2) The expression “laughing" is also restricted. Conversely, the results from FD show greater prompt consistency. These findings indicate that current identity embedding methods may not effectively address “the balance between identity fidelity and prompt consistency".
@@ 3.2 ID Embedding’s Effect on Attention Layers
Figure 2: Analysis on cross-attention maps of facial personalization models. Left: token-wise attention map visualization. Right: interpolation experiment on PD and FD’s cross-attention maps.
OBSERVATION 1: ID embeddings disrupt cross-attention layers. To investigate the underlying causes of this phenomenon, we conduct a visualization experiment inspired by  hertz2023prompttoprompt ; alaluf2024cross ; tewel2024training , because identity embedding is primarily fused into cross-attention layers. Columns (b-d) of Fig.2’s left part present the visualization results of cross-attention maps for both PD and FD processes, focusing on identity embeddings and the attribute-specific token embeddings for “black curly hair" and “laughing". It reveals that the identity embedding significantly amplifies activation in facial regions while simultaneously reducing activation for other tokens and disrupting their typical interactions within the cross-attention mechanism. Additionally, we observe that during the FD process, prompt consistency remains preserved. This observation underscores an insight: although most personalization models are well fine-tuned, their intrinsic capability to generate faces with high prompt consistency is still preserved but unexpectedly overridden by external identity embeddings. Notably, this latent capability can be effectively activated via the FD procedure, as its name “Foundation Denoising" indicates.
OBSERVATION 2: Personalized cross-attention layers are highly susceptible. We conduct another experiment on cross-attention layers, which involves incorporating a portion of the cross-attention maps from the FD process into those of the PD process. The formula for this approach is
@@
where $A^{p}_{:,:,m}$ represents the cross-attention map of PD’s identity embedding, and $A^{f}_{:,:,n}$ represents the cross-attention map of the FD’s zero-valued embedding. The parameter $\alpha$ controls the weight of the FD’s cross-attention map that is injected. The results are shown on the right part of Fig.2. It is observed that, as the weight of the map from the FD process increases, the model quickly loses identity fidelity, even though it regains the facial attributes that were missing before. This experiment demonstrates that cross-attention layers are rather susceptible. To preserve the models’ capacity for identity preservation, it is better to leave these well-trained cross-attention modules unaltered.
Our preliminary findings demonstrate that identity embeddings in cross-attention layers are disrupting personalization models’ prompt consistency. In contrast, its strong ability of identity extraction makes it rather challenging to be modified. Given this “dilemma", we aim to explore a new approach from the perspective of self-attention, inspired by  gu2024photoswap ; tewel2023key . Since most personalization models introduce minimal modifications to self-attention layers, it is reasonable to assume that the aforementioned hidden foundation knowledge is preserved within them. By enhancing the self-attention layers in personalization models while keeping cross-attention layers intact, we anticipate achieving better alignment of facial attributes in personalized generation.
@@ 4 Methodology
Building upon our investigation in Sec.3, we propose FreeCure, a training-free framework designed to improve the prompt consistency of facial personalization models (see Fig.3). Initially, we develop a foundation-aware self-attention (FASA) mechanism to integrate localized attributes from the FD process into the PD process (Sec.4.1). Subsequently, we apply an asymmetric prompt guidance (APG) strategy to reconstruct more abstract attributes (Sec.4.2).
Figure 3: Overview of FreeCure. For a personalization model $\epsilon_{\theta}$, we first introduce (a): dual inference paradigm to generate faces with/without identity ($I_{p}$ and $I_{f}$), where $I_{f}$ without identity embedding shows better prompt consistency. Next, we leverage a segmentation model $\Psi(\cdot)$ to derive related masks of target attributes with clear spatial information (hair, sunglasses, etc.) and merge them into a mask $\mathcal{M}$. In (b): we modify the original self-attention modules with our proposed FASA (c), which concatenates key and value matrices of FD process and PD process, together with a scaling mask to achieve the attribute injection. Finally, we utilize a simple yet effective strategy (d): asymmetric prompt guidance (APG) to restore abstract attributes (e.g., expressions).
@@ 4.1 Foundation-Aware Self-Attention
Obtain Masks of Spatial Localized Attributes. As shown in Fig.3 (a), we adapt the dual inference branches which are identical to the FD/PD process introduced in Sec.3 for personalization models to generate the personalized face $I_{p}$ with unsatisfactory prompt consistency as well as the foundation face $I_{f}$ with high prompt consistency. Next, we utilize an external face parsing model, denoted as $\Psi(\cdot)$, to extract the binary mask $M_{i}$ corresponding to $A_{i}$ from the foundation outputs $I_{f}$. Generally, when restoring multiple attributes, we simply need to compute the different masks respectively and merge them $\mathcal{M}=\bigcup\{M_{i}\}$. $\mathcal{M}$ contains the spatial information of attributes that align with the target prompt, which will play an important role in the attribute restoration process that will be mentioned in the next part.
FASA Mechanism. FASA is the core module that links the information between FD and PD processes and enables personalization models to restore the attributes via their foundation knowledge. As shown in Fig.3(c), we at first identify the self-attention modules in the model. Specifically, in each timestep $t$ and attention layer $l$, we denote PD and FD’s key, query, and value matrices as $\mathcal{KQV}^{tl}_{p}=\{K^{tl}_{p},Q^{tl}_{p},V^{tl}_{p}\}$ and $\mathcal{KQV}^{tl}_{f}=\{K^{tl}_{f},Q^{tl}_{f},V^{tl}_{f}\}$. Second, we concatenate the key and value matrices of FD process to those of PD process: $\hat{K^{tl}}=[K^{tl}_{p},K^{tl}_{f}]$, $\hat{V^{tl}}=[V^{tl}_{p},V^{tl}_{f}]$. Thus, by omitting the labels of $t$ and $l$ for simplicity, the operation of FASA is
@@
Fine-grained Restoration with Scaling Masks. The approach described in Eq.3 allows the PD process to obtain information from the FD process. However, we observe that the method also retains a substantial degree of unrelated features, leading to a notable loss of identity fidelity, as shown in Fig.4. To constrain attribute restoration in a specific region without disrupting identity information, we apply pre-calculated masks $\mathcal{M}$ of different attributes to the similarity map. Additionally, to further enhance FASA’s performance, we introduce an additional scaling factor, denoted as $\omega$, to control the magnitude of injecting attribute features from FD to PD. Therefore, the enhanced FASA mechanism can be written as
@@
Where $\mathbf{1}$ denotes a matrix with all elements equal to 1, designed to preserve attention to the original features within the PD process. $\odot$ represents the Hadamard product.
In the full-attention layers of Diffusion Transformers (DiT) such as FLUX, visual and textual information is integrated into a unified sequence representation $[X;C]$. Within this architecture, the FASA mechanism operates in a similar manner, with a key distinction: the attribute mask is applied solely to the component derived from the visual queries of the PD branch ($Q_{p}^{X}$) and the visual keys of the FD branch ($K_{f}^{X}$). This selective masking strategy preserves the original cross-modal attention patterns between visual and textual elements, resembling OminiControl tan2024ominicontrol :
$\mbox{FASA}_{flux}(\mathcal{KQV}_{p},\mathcal{KQV}_{f})=\mbox{Softmax}(\frac{\mathcal{M}(\omega)_{flux}\odot Q_{p}\hat{K}^{T}}{\sqrt{d}})\hat{V}$, (5)
$\mathcal{M}(\omega)_{flux}=\begin{pmatrix}\mathbf{1}_{l_{1}\times l_{1}}&\mathbf{1}_{l_{1}\times l_{2}}&\omega\mathcal{M}_{l_{1}\times l_{1}}\\ \mathbf{1}_{l_{2}\times l_{1}}&\mathbf{1}_{l_{2}\times l_{2}}&\mathbf{0}_{l_{2}\times l_{1}}\end{pmatrix}$ (6)
Where $Q_{p}=[Q_{p}^{X};Q_{p}^{C}]$ is the original query matrix of PD branch, and $\hat{K}=[K_{p}^{X};K_{p}^{C};K_{f}^{X}],\hat{V}=[V_{p}^{X};V_{p}^{C};V_{f}^{X}]$ are FASA’s enhanced key and value matrices which integrates information from both PD and FD branches. With such enhancement, attributes of the FD process can be successfully extracted and precisely injected into the PD process without disrupting the personalization model’s ability of maintaining identity fidelity.